OpenStudy (anonymous):

how do I leave 8 to the 11th power/ 8 to the 5th power in exponential notation? or how do I write it in that form?

OpenStudy (anonymous):

\[a^m/a^n = a ^{m-n}\]

OpenStudy (anonymous):

Can you write it in exponential form now ?

OpenStudy (anonymous):

no i dont get it

OpenStudy (anonymous):

Did you understand my first post?

OpenStudy (anonymous):

8 11-5

OpenStudy (anonymous):

or 8 to the 6th power

OpenStudy (anonymous):

Yes 8^11-5 = 8^6

OpenStudy (anonymous):

if i leave it in exponential notation would it be 8 to the 6th or?

OpenStudy (anonymous):

Yes it would be 8^6

OpenStudy (anonymous):

thank you so much...

OpenStudy (anonymous):

ur welcome :D